7. It takes a plane 3 hours to make a 900 mile trip with the wind. The return trip against the same wind takes 3 hours and 24 minutes. Find the speed of the plane in still air and the speed of the wind.

Answer by richwmiller(17219)

r*t=d
with the wind
(r+w)*3=900
(r+w)=900/3
r+w=300
against the wind
(r-w)*3.4=900
(r-w)=900/3.4
r-w=264.705882
r+w=300
add
r+r=300+264.705882
2r=564.705882
r=282.352941 mph rate
282.352941-w=264.705882
w=282.352941-264.705882
w=17.6470588 mph wind
check
(282.352941+17.6470588)*3=900
300.0*3=900
900.0=900
ok
(r-w)*3.4=900
(282.352941-17.6470588)*3.4=900
264.705882*3.4=900
900.0=900
ok